OpenStudy (anonymous):

Solve 2x + y = 7 x – y = –4 a. (1, 5) b. (–1, 5) c. (1, –5) d. (–1, –5)

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

2x+y=7 y=7-2x out this y value in x-y=-4 and solve for x x-7+2x=-4 3x=-4+7 x=3 plug this x value in x-y=-4 1-y=-4 -y=-4-1 -y=-5 y=5 (1,5)

OpenStudy (anonymous):

Thanks you very much for your help :)

OpenStudy (anonymous):

welcome!!